# MetaLint: Generalizable Idiomatic Code Quality Analysis Through Instruction-Following and Easy-to-Hard Generalization

### v2

@@ 5.2 Evaluating Easy-to-Hard Generalizations
Table 3: Easy-to-Hard Generalization on PEP Idioms: We evaluate the effect of different MetaLint training setups (SFT, RS-SFT, and RS-DPO) on Qwen3-4B (with and without reasoning) and Llama3.2-3B. Models are trained on easy synthetic Python Ruff idioms and tested on hard manually curated PEP idiom detection data which can't be handled by linters or static analyzers (section 4.3). Best score across the compared training setups per model are bolded.
5.3 Benchmarking on Hard Idioms
Detection: In terms of detection F-score, the non-CoT MetaLint model is competitive with o3-mini and GPT-5 but is outperformed by some larger open-source models (e.g., Qwen3-32B with CoT, DeepSeek-R1-Distill-Qwen-32B with CoT, and GPT-oss-120B) and closed-source models (GPT-4o, GPT-4.1, and o4-mini). However, the non-CoT model achieves the highest detection recall among all evaluated models, while the CoT model ranks among the top in precision, surpassed only by Qwen3-32B with CoT and o4-mini.
Localization: For localization, the MetaLint models lag behind larger 32B and 120B models (such as Qwen3-32B, Qwen2.5Coder-32B, and DeepSeek-R1-Distill-Qwen-32B) and the GPT models, but perform comparably to o3-mini (statistical significance analysis in Appendix E.2) and outperform GPT-oss-20B. This is notable given that the MetaLint models are much smaller (4B parameters), trained only on synthetic data derived from easy idioms, and that the non-CoT model does not use test-time compute.
Overall, the strong results, especially the best-in-class recall of the non-CoT model, demonstrate the effectiveness of our framework in achieving easy-to-hard generalization. This is enabled by training on synthetic data with easy idioms and by encouraging adaptive reasoning through instruction fine-tuning and DPO rather than relying on rote memorization.
@@ 5.3 Benchmarking on Hard Idioms
Table 4: Benchmarking on Hard Idioms: Results comparing state of the art code and reasoning models on the hard PEP benchmark to contextualize the gains achieved with MetaLint training. The best scores are bolded and second best and underlined.
@@ 6 Conclusion and Future Work
Our results show that MetaLint training fosters adaptive reasoning over idiom specifications rather than rote memorization. We observe generalization to unseen idioms in Python and Java, across three linters (Ruff, PMD, JEP tree-sitter), two model families (Qwen, Llama), reasoning and non-reasoning settings, and multiple scales (3B, 4B, 8B). Easy-to-hard generalization occurs from linter-detectable Ruff idioms to harder PEP idioms, with SFT teaching output formatting and DPO enabling true generalization. Compared to state-of-the-art code and reasoning models, MetaLint-trained Qwen models have detection comparable with o3-mini and GPT-5, achieving highest recall (non-CoT) and third-best precision (CoT). Localization lags but surpasses GPT-oss-20B with only 4B parameters and no test-time compute and is comparable to o3-mini, demonstrating efficiency. These results highlight the effectiveness of instruction fine-tuning and preference optimization on synthetic data for reasoning and generalization, even with scarce annotated examples. For mechanically easy idioms, linters remain cost-effective, but MetaLint enables detection of abstract idioms, supporting personalized, evolving code quality standards. We plan to release code and data for reproducibility. Future work includes training for automated refactoring and exploring advanced RL methods like Group Relative Policy Optimization (GRPO) Shao et al. (2024).
Acknowledgments
@@
Appendix A Limitations
Despite the promising results achieved by MetaLint, our work has some limitations that we plan to address in future research. For the CoT setting, we didn't explore whether non-CoT models can be trained to effectively produce CoT-style reasoning with supervision from a teacher model. We also explored self-improvement strategies for RS-SFT data generation in cases where the base model failed, such as STaR (Zelikman et al., 2022), but found it challenging to generate CoTs that do not directly reference provided hints, which risks contaminating the training data. As a result, we adopted a simpler rejection sampling or RS-SFT strategy. Furthermore, our approach does not yet incorporate more advanced reinforcement learning techniques such as Group Relative Policy Optimization (GRPO) (Shao et al., 2024) using our verifiable linter-based reward model, or curriculum learning methods to control the progression of idiom difficulty within synthetic training data. Our current experiments also focus on training on one language at a time, such as only Python or Java. Future work will explore joint training and extension to more programming languages like JavaScript, Ruby, Go, etc., as well as cross-language generalization by training on Python idioms and evaluating on Java idioms, and vice versa. Finally, while we do not evaluate or train for refactoring of the idiom-violating code, we plan to do so in future work.
Appendix B More Related Work
Easy-to-Hard Generalization. Research shows that training on simpler problems enhances generalization to harder ones in math, algorithms, and code, motivating its application to code quality analysis. In math reasoning, models trained on easier problems (e.g., level 1–3) consistently generalize better to harder benchmarks (e.g., level 4–5) (Bai et al., 2024; Shafayat et al., 2025; Parashar et al., 2025). Several works emphasize the importance of selecting high-quality supervision for harder problems (He et al., 2024). Beyond math, Sun et al. (2024a) shows that reward models trained on simple code and math problems improve performance on complex ones. Broader studies on multi-task and length generalization (Hu et al., 2025) and differentiable programming (Gaunt et al., 2016) reveal how structural simplicity during training can lead to robustness on longer or more complex reasoning instances, including code. Zhang et al. (2024a) reinforces this by evaluating reward models on algorithmic tasks like string manipulation and demonstrating transfer from simpler to harder formats. Drawing inspiration from this work, we train MetaLint on large-scale synthetic data covering easily detectable code idioms handled by rule-based linters, and hypothesize that these simple patterns serve as stepping stones toward generalizing to complex, novel PEP idioms.
Appendix C Method Additional Details
C.1 MetaLint Instruction Following Prompt
@@
We used the following instruction following style prompt to train the model with synthetic Ruff idiom data for the meta-linting task:
@@
An example input with the code file and idiom spec populated as well as the expected JSON style output is shown below:
@@
C.2 DPO Contrastive Pair and RS-SFT Sampling Details
@@
For RS-DPO sampling (in both CoT and non-CoT settings), we used the standard MetaLint instruction-following prompt with the SFT models. In contrast, for RS-SFT output sampling from the untrained model, we employed the expanded ``Baseline Inference Prompt'' described in Section C.4.
C.3 Training Hyperparameters and Computational Environment
@@
C.4 Baseline Inference Details
@@
Additionally, we use an expanded prompt (Baseline Inference Prompt) compared to the one used for MetaLint, specifically adding more details about output formatting to ensure all baselines have a fair chance and do not suffer performance drops due to formatting mismatches. For the same reason, we also allow certain relaxations in output formatting during evaluation on the PEP Hard Idiom Benchmark.
@@
Table 5: JEP Idiom Specifications (1/3): This table presents 15 idioms across 8 JEPs, including both ``before'' (old best practice) and ``after'' (updated best practice) patterns. The JEP# column lists the JEP number, the JEP title specifies the idiom topic, and the parenthesized value indicates whether it is a before or after pattern. The Definition, Example, and Tree-Sitter Queries columns provide the idiom definition, minimal Java examples shown to the LLM as instructions, and the queries used to flag idioms for synthetic data creation.
@@
Table 6: JEP Idiom Specifications (2/3)
@@
Table 7: JEP Idiom Specifications (3/3)
C.5 PMD Idiom Specifications
We scrape PMD idioms specification from the Java section of the PMD rules documentationhttps://docs.pmd-code.org/latest/pmd_rules_java.html. The PMD instructions are more complex and more ambiguous than our handcrafted JEP specifications because the examples are more verbose and don't pinpoint the specific lines that should be flagged as idiom violations, as can be seen in the example below.
@@ C.5 PMD Idiom Specifications
Table 8: List of JEPs addressed by our tree-sitter synthetic data. The JEP# and Title column indicate the number and title of the JEP while JDK# and Release Date indicate the JDK needed for compilation to be able to use the JEP features. The Before and After columns indicate whether we include rules/patterns to flag the old idiom or new idiom introduced by the JEP.
@@
Appendix D Additional Experimental Details
D.1 Evaluation Metrics
Let $I$ denote an idiom, $M_{I}$ its corresponding meta task specification, $f\in\mathcal{F}$ a code file, $V_{f,I}$ the ground truth set of violating line numbers, and $\hat{y}=V^{\Phi}_{f,I}$ the model predicted violations. For each dataset instance with input prompt $x$ and ground truth set of line numbers $y$, $(x,y)=(\{f,M_{I}\},V_{f,I})\in\mathcal{D}$.
@@ D.1 Evaluation Metrics
D.2 Idioms Chosen for Ruff Idiom Transfer Dataset
Table 9 lists the Ruff idioms used in the SFT training and synthetic transfer evaluation test sets. Idioms are grouped by their source linter and cover a range of syntax, semantics, naming, and upgrade-related rules.
@@ D.2 Idioms Chosen for Ruff Idiom Transfer Dataset
Table 9: Ruff idioms included in the supervised training and transfer evaluation test sets. Test set idioms span both overlapping linters and novel ones not seen during training.
@@
D.3 DPO No Violation Fraction Ablations
@@
Fraction of
NV data
Detection Localization
@@
Table 10: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Qwen3-4B model without CoT. Including 0% yields the highest recall and best line-level localization but reduces precision due to more false positives and lower accuracy in predicting NO VIOLATIONS FOUND. Conversely, including 100% improves precision but leads to reduced recall and localization performance. All rows report the performance at the best training step, selected based on a balance of detection and localization F-score on the Ruff Idiom Transfer test set.
Fraction of
NV data
Detection Localization
@@
Fraction of
NV data
Detection Localization
@@
D.4 PEP Benchmark Creation Additional Details
As discussed in section 4.3 we use some high recall heuristics to find promising candidates for detecting the selected hard PEP idioms. These are summarized in Table 13, 14 and 15.

### v3

@@ 5.2 Evaluating Easy-to-Hard Generalization
Table 4: Benchmarking on Hard Best Practices: Results comparing state of the art code and reasoning models on the hard PEP benchmark to contextualize the gains achieved with MetaLint training. The best scores are bolded and second best and underlined.
@@
5.3 Benchmarking on Hard Best Practices
Detection: For detection F-score, the non-CoT MetaLint model is competitive with o3-mini and GPT-5 but is outperformed by some larger open-source models (e.g., Qwen3-32B with CoT, DeepSeek-R1-Distill-Qwen-32B with CoT, and GPT-oss-120B) and closed-source models (GPT-4o, GPT-4.1, and o4-mini). However, the non-CoT model achieves the highest detection recall among all evaluated models, while the CoT model ranks among the top in precision, surpassed only by Qwen3-32B with CoT and o4-mini.
Localization: For localization, the MetaLint models lag behind larger 32B and 120B models (such as Qwen3-32B, Qwen2.5Coder-32B, and DeepSeek-R1-Distill-Qwen-32B) and the GPT models, but perform comparably to o3-mini (statistical significance analysis in Appendix G.2) and outperform GPT-oss-20B. This is notable given that the MetaLint models are much smaller (4B parameters), trained only on synthetic data derived from easy best practices, and that the non-CoT model does not use test-time compute. We also identify localization errors and their frequency in Table 32.
Overall, our models perform comparably to state-of-the-art systems, with the non-CoT Qwen3-4B variant achieving the highest recall, and both MetaLint-trained Qwen and LLaMA models exhibiting evidence of easy-to-hard generalization. This aligns with Chang & Bisk (2025), which highlights the role of reorganizing supervision and learning paradigm, rather than scaling or optimization alone, in enabling generalization along difficulty progressions.
@@ 6 Conclusion and Future Work
Our results show that MetaLint encourages inductive transfer over best practice specifications, rather than memorization of fixed rules. We observe generalization to unseen best practices across Python and Java, multiple linters, model families, reasoning settings, and model scales. Models trained with MetaLint also exhibit easy-to-hard generalization, transferring from linter-detectable best practices to harder PEP violations. With only 4B parameters, MetaLint-trained Qwen models achieve detection performance comparable to strong code and reasoning systems, achieving the highest recall and competitive precision. Localization remains more challenging but is competitive with o3-mini and surpasses GPT-oss-20B without test-time compute. Overall, these results support the view that reorganizing supervision, rather than scaling alone, enables inductive generalization beyond the training distribution (Chang & Bisk, 2025). MetaLint is not a replacement for static analysis, but a proof of concept that inductive learning can improve LLM performance along a natural difficulty progression in code quality analysis. Future work will target automated refactoring and better RL training like Group Relative Policy Optimization (GRPO) (Shao et al., 2024).
Acknowledgements
@@
Impact Statement
This work aims to improve the reliability and generalization of automated code quality analysis by enabling models to adapt to evolving and previously unseen best practices. By reducing reliance on brittle, static rule sets, MetaLint has the potential to lower maintenance costs for software tooling and improve developer productivity and security outcomes, particularly in settings where best practices change rapidly. However, as with any automated code analysis system, incorrect or overconfident recommendations could mislead developers or create false assurance if deployed without human oversight. We mitigate this risk by grounding supervision in existing static analysis signals and framing outputs as assistive diagnostics rather than authoritative judgments. We do not anticipate significant negative societal impacts beyond those common to developer tooling, and we expect the primary effects of this work to be beneficial.
Appendix A Appendix
We organize the Appendix in five parts: 1) Limitations of our work (Appendix B), 2) Clarifications and definitions of terms used in the paper (Appendix C), 3) Additional related work about instruction following and generalization (Appendix D), 4) Method details like prompts and hyperparameters (Appendix E), 5) Additiomal experimental details (Appendix F) and 6) Additional results and ablations (Appendix G).
Appendix B Limitations
Despite promising results, including evidence of easy-to-hard generalization, MetaLint has several limitations that suggest directions for future work:
  • Chain-of-Thought (CoT) reasoning: We did not explore whether non-CoT models can be trained to generate CoT-style reasoning using supervision from a teacher model.
  • Self-improvement and data generation: We attempted self-improvement strategies for RS-SFT data generation (e.g., STaR (Zelikman et al., 2022)) when the base model failed. However, generating CoTs that do not directly reference provided hints proved challenging and risks contaminating the training data. We therefore adopted a simpler rejection sampling (RS-SFT) strategy.
  • Preference optimization methods: We only experimented with DPO for preference optimization. Alternative approaches, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017) or Group Relative Policy Optimization (GRPO) (Shao et al., 2024), combined with our verifiable linter-based reward model, could be explored.
  • Difficulty progression and curriculum: Our current setup does not implement a fine-grained difficulty progression or a curriculum over meta-task abstractions. Incorporating such structure could encourage more systematic induction over code idioms and code quality concepts.
  • Multilingual training: Experiments focus on one programming language at a time (Python or Java). Future work could explore joint training across multiple languages, cross-language transfer, and structured learning that encourages abstraction across similar language constructs (e.g., Python’s PEP 557: Data Classes vs. Java’s JEP 395: Records).
  • Task scope: We focus exclusively on best practice violation detection and do not yet explore inductive transfer over code refactoring or other code quality tasks.
Appendix C Clarification and Term Definitions
Because this paper studies a software engineering–heavy problem setting, we clarify the key technical terms used throughout.
Code Idiom / Programming Idiom.
A code idiom is a recurring, language-specific pattern of code that conveys a particular semantic intent or programming practice beyond what is expressed by its literal syntax.
Code Quality Analysis.
Code quality analysis refers to the systematic evaluation of source code with respect to software engineering attributes such as readability, complexity, maintainability, and adherence to best practices, using automated tools or model-based methods. Although best practice analysis is only one component of code quality analysis, we use these terms interchangeably in this paper for brevity.
Best Practice Violation Detection.
Best practice violation detection is a subproblem of code quality analysis that involves identifying and localizing code fragments that violate established or recommended programming practices. In this work, we operationalize best practice violation detection as the task of identifying problematic code idioms, i.e., recurring patterns that correspond to non-recommended practices. While this formulation is more restrictive than the full space of best practices, it enables a systematic task structure that supports controlled supervision and generalization. Additionally we often use the shorthand best practice idiom or idiom to refer to the problematic idiom that needs to be detected to flag a best practice violation.
Appendix D More Related Work
Instruction Following and Generalization. Instruction tuning has emerged as a powerful form of meta-learning, enabling cross-task generalization by training models to interpret and follow natural language instructions rather than memorizing fixed tasks. Prior work shows that exposure to diverse task instructions allows models to extract underlying task abstractions and apply them to unseen settings (Mishra et al., 2021b; Wang et al., 2022). Large-scale instruction tuning further improves zero- and few-shot generalization across tasks and modalities (Wei et al., 2021; Chung et al., 2022; Gao et al., 2021; Iyer et al., 2022; Brown et al., 2020). Instructions serve as high-density task representations, substituting for explicit supervision (Puri et al., 2022) and enabling generalization even with minimal labeled data or pseudo-labeled examples (Gu et al., 2022). Studies also show that instruction diversity drives generalization: varied instructions outperform repeated exposure to identical formats (Charton et al., 2024). This effect holds across domains, including program synthesis, where task-level prompting facilitates generalization in code generation models (Niu et al., 2023). SELF-GUIDE (Zhao et al., 2024) performs task-specific instruction following using synthetic data, demonstrating effectiveness but relying entirely on LLM-generated data without verifiers. These results suggest that instruction tuning functions as task-level meta-learning, enabling models to adapt to new tasks through natural language. Building on this insight, we model individual code quality idioms as separate meta-tasks and generate large-scale synthetic data for each task to support cross-idiom generalization. This approach allows the trained model to adapt to new idioms and evolving best practices.
Appendix E Method Additional Details
E.1 MetaLint Instruction Following Prompt
@@
We used the following instruction following style prompt to train the model with synthetic Ruff best practice idiom data for the meta-linting task:
@@
An example input with the code file and best practice idiom spec populated as well as the expected JSON style output is shown below:
@@
E.2 DPO Contrastive Pair and RS-SFT Sampling Details
@@
For RS-DPO sampling (in both CoT and non-CoT settings), we used the standard MetaLint instruction-following prompt with the SFT models. In contrast, for RS-SFT output sampling from the untrained model, we employed the expanded ``Baseline Inference Prompt'' described in Section E.4.
E.3 Training Hyperparameters and Computational Environment
@@
E.4 Baseline Inference Details
@@
Additionally, we use an expanded prompt (Baseline Inference Prompt) compared to the one used for MetaLint, specifically adding more details about output formatting to ensure all baselines have a fair chance and do not suffer performance drops due to formatting mismatches. For the same reason, we also allow certain relaxations in output formatting during evaluation on the PEP Hard Best Practice Benchmark.
@@
Table 5: JEP Best Practice Idiom Specifications (1/3): This table presents 15 best practice idioms across 8 JEPs, including both ``before'' (old best practice) and ``after'' (updated best practice) patterns. The JEP# column lists the JEP number, the JEP title specifies the idiom topic, and the parenthesized value indicates whether it is a before or after pattern. The Definition, Example, and Tree-Sitter Queries columns provide the idiom definition, minimal Java examples shown to the LLM as instructions, and the queries used to flag problematic idioms for synthetic data creation.
@@
Table 6: JEP Best Practice Idiom Specifications (2/3)
@@
Table 7: JEP Best Practice Idiom Specifications (3/3)
@@
E.5 PMD Best Practice Idiom Specifications
We scrape PMD idioms specification from the Java section of the PMD rules documentationhttps://docs.pmd-code.org/latest/pmd_rules_java.html. The PMD instructions are more complex and more ambiguous than our handcrafted JEP specifications because the examples are more verbose and don't pinpoint the specific lines that should be flagged as best practice idiom violations, as can be seen in the example below.
@@ E.5 PMD Best Practice Idiom Specifications
Table 8: List of JEPs addressed by our tree-sitter synthetic data. The JEP# and Title column indicate the number and title of the JEP while JDK# and Release Date indicate the JDK needed for compilation to be able to use the JEP features. The Before and After columns indicate whether we include rules/patterns to flag the old problematic idiom or new recommended idiom introduced by the JEP.
@@
Appendix F Additional Experimental Details
F.1 Evaluation Metrics
Let $I$ denote an best practice idiom, $M_{I}$ its corresponding meta task specification, $f\in\mathcal{F}$ a code file, $V_{f,I}$ the ground truth set of violating line numbers, and $\hat{y}=V^{\Phi}_{f,I}$ the model predicted violations. For each dataset instance with input prompt $x$ and ground truth set of line numbers $y$, $(x,y)=(\{f,M_{I}\},V_{f,I})\in\mathcal{D}$.
@@ F.1 Evaluation Metrics
F.2 Idioms Chosen for Ruff Best Practice Transfer Dataset
Table 9 lists the Ruff best practice idioms used in the SFT training and synthetic transfer evaluation test sets. Idioms are grouped by their source linter and cover a range of syntax, semantics, naming, and upgrade-related rules.
Table 9: Ruff best practice idioms included in the supervised training and transfer evaluation test sets. Test set idioms span both overlapping linters and novel ones not seen during training.
@@ F.2 Idioms Chosen for Ruff Best Practice Transfer Dataset
F.3 DPO No Violation Fraction Ablations
@@
Table 10: Effect of varying the fraction of NO VIOLATIONS FOUND instances in the training data for MetaLint Qwen3-4B model without CoT. Including 0% yields the highest recall and best line-level localization but reduces precision due to more false positives and lower accuracy in predicting NO VIOLATIONS FOUND. Conversely, including 100% improves precision but leads to reduced recall and localization performance. All rows report the performance at the best training step, selected based on a balance of detection and localization F-score on the Ruff Best Practice Transfer test set.
Fraction of NV data Detection Localization
@@
Fraction of NV data Detection Localization
@@
Fraction of NV data Detection Localization
@@
Table 13: Effect of Simple Ensembling Strategies on Detection and Localization Performance: Comparison of intersection and union-based ensembling of SFT and RS-SFT models on Qwen3-4B. The intersection strategy retains only line numbers predicted by both models, while the union strategy aggregates all predicted line numbers from either model. Both ensembling approaches underperform the individual models, with the intersection strategy exhibiting particularly poor precision and recall, suggesting that the observed precision–recall differences are driven by training setup rather than a fundamental trade-off.
Ensembling Strategy Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B DPO Intersection (non-CoT & CoT) 0.0667 0.0033 0.0063 0.0000 0.0000 0.0000
Qwen3-4B DPO Union (non-CoT & CoT) 0.6188 0.4187 0.4995 0.1952 0.1251 0.1524
Table 14: Effect of Removing NO VIOLATIONS (NV) Data on Hard PEP Benchmark Performance: Full detection and localization results for Qwen3-4B trained with SFT+RS-DPO under different NV data settings. The 0% NV variant achieves substantially higher recall than the model reported in the main paper, with only a moderate reduction in precision. Since the precision drop is smaller than the recall gain, this leads to an improved overall F score, confirming that the originally reported recall was an underestimate and further strengthening our state-of-the-art recall claim.
Model Detection Localization
$P_{Det}$ $R_{Det}$ $F_{Det}$ $P_{Loc}$ $R_{Loc}$ $F_{Loc}$
Qwen3-4B (SFT + RS-DPO) 0.7031 0.7043 0.7037 0.3536 0.1930 0.2497
Qwen3-4B (SFT + RS-DPO, 0% NV) 0.6658 0.8037 0.7283 0.3797 0.2063 0.2674
Qwen3-4B w CoT (RS-SFT + RS-DPO) 0.9303 0.4958 0.6468 0.3482 0.2169 0.2673
F.4 PEP Benchmark Creation Additional Details
As discussed in section 4.3 we use some high recall heuristics to find promising candidates for detecting the selected hard PEP best practice idioms. These are summarized in Table 15, 16 and 17.